How many liters (L) of a 20% solution must be mixed with 40 L of a 80% solution to get a 50% solution?
My teacher wants us to make 2 equations. We solve them by plugging them into the calculator. I just need to know how to get 2 equations out of this. Thank you so much!

Answer by stanbon(75887)

How many liters (L) of a 20% solution must be mixed with 40 L of a 80% solution to get a 50% solution?
-----
Quantity Equation: x + 40 = y
Ingrediant Eq::: 0.20x + 0.80*40 = 0.50y
----------------------------------------------
Rearrange and multiply the 1st equation by 20
Rearrange and multiply the 2nd equation by 100
----------------------------------------------
20x - 20y = -800
20x - 50y = -3200
-------------------
Cheers,
Stan H.